If cot=1.5 and it is in quadrant 3 whats the value of sin
Answer by stanbon(75887)

If cot=1.5 and it is in quadrant 3 whats the value of sin
----
sin is negative in the 3rd quadrant.
------
cot = x/y = 3/2
---
I x = 3 and y = 2, then r = sqrt[2^2+3^2] = sqrt(13)
---------
Answer: sin = y/r = -2/sqrt(13)
------------
Cheers,
Stan H.